OpenStudy (anonymous):

Discs of radius 4cm are cut from a rectangular sheet of length 84cm and width 24cm. How many complete discs can be cut out? what is the total area of the disks cut? what is the area of the sheet wasted?

OpenStudy (anonymous):

nika first figure out how many discs you can make from the sheet

OpenStudy (anonymous):

if your radius is 4, your diameter is 2r so D=8 next figure out how many rows and columns you can make

OpenStudy (anonymous):

first take your length of the sheet and find how many rows you can make of circles... next find howm many diameters the width can have multiply these numbers together this will tell you how many discs you can make next use the area formula of a circle and multiply it by the amount of discs you can make

OpenStudy (anonymous):

I'll give you a hint solve these 8x=84 8x=24

OpenStudy (anonymous):

solve x for both of these really you should make one of them a y variable.

OpenStudy (anonymous):

okay i figured out that part i think, i got 10.5 and 3

OpenStudy (anonymous):

D = 8cm 8x = 24 8x = 84 Since length is 84cm, about 10.5 discs can be made horizontally. We will not count the 0.5 since that is invalid for this situation. Now, since the width is 24cm, 3 discs can be made vertically. Now we multiply 10 x 3 = 30 discs which can be made in total. A=10(3.14r^2) A=10(3.14x4^2) A=502.4cm^2 now, since 0.5 disc is invalid, that will be considered as waste. 0.5 x 3 = 1.5 discs are wasted A = 1.5(3.14(4)^2) A = 75.36cm^2